ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 239, 2020

Alvarenga, R. Hall algebras and graphs of Hecke operators for elliptic curves

215

Bárány, B., A. Käenmäki and I. D. Morris Domination, almost additivity, and thermodynamic formalism for planar matrix cocycles

173

Burness, T. C. and S. Harper Finite groups, 2-generation and the uniform domination number

271

Czudek, K. and T. Szarek Ergodicity and central limit theorem for random interval homeomorphisms

75

Eleftheriou, P. E., A. Günaydin and P. Hieronymi Structure theorems in tame expansions of o-minimal structures by a dense set

435

Fox, J., J. Pach and A. Suk The Schur–Erdős problem for semialgebraic colorings

39

Futorny, V., D. Grantcharov, L. E. Ramirez and P. Zadunaisky Gelfand–Tsetlin theory for rational Galois algebras

99

Giambruno, A. and D. La Mattina Codimensions of star-algebras and low exponential growth

1

Grantcharov, D. See V. Futorny Günaydin, A. See P. E. Eleftheriou Harper, S. See T. C. Burness Hieronymi, P. See P. E. Eleftheriou Kabluchko, Z., J. Prochno and C. Thäle Intersection of unit

balls in classical matrix ensembles

129

Käenmäki, A. See B. Bárány Kislyakov, S. V. and I. K. Zlotnikov Interpolation for

intersections of Hardy-type spaces

21

Levit, A. On Benjamini–Schramm limits of congruence subgroups  59

La Mattina, D. See A. Giambruno Morris, I. D. See B. Bárány Pach, J. See J. Fox Prochno, J. See Z. Kabluchko Ramirez, L. E. See V. Futorny Sheth, M. Locally analytic representations in the étale coverings of

the Lubin–Tate moduli space

369

Suk, A. See J. Fox Szarek, T. See K. Czudek Thäle, C. See Z. Kabluchko Zadunaisky, P. See V. Futorny Zlotnikov, I. K. See S. V. Kislyakov